OpenStudy (anonymous):

Simplify square root 2475

OpenStudy (anonymous):

A) \[9\sqrt{30}\] B)\[11\sqrt{15}\] C)\[15\sqrt{11}\] D)\[45\sqrt{11}\]

OpenStudy (anonymous):

15sqrt(11)

OpenStudy (anonymous):

try it this way: \[\sqrt{2475}=\sqrt{5 \times 495}=\sqrt{5 \times 5 \times 99} = \sqrt{5 \times 5 \times 3 \times 3 \times 11} = 5 \times 3 \times \sqrt{11}=15\sqrt{11}\]